Circular No. 6131

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

NOVA CIRCINI 1995 A. C. Gilmore and P. M. Kilmartin report the following precise position

of N Cir 1995 from a film obtained with the 0.15-m f/15 astrograph at Mt. John University Observatory on Jan. 29.55 UT: α = 14^{h}44^{m}53^{s}.48, δ = −63^{o}53′55′′.6 (equinox 2000.0; rms residual from eight PPM reference stars was 0′′.1 in α and 0′′.4 in δ). Further photometry (Mt. John 0.6-m reflector): Jan. 29.448 UT, V = 7.92, U − B = −0.44, B − V = +0.37, V − R = +1.52, V − I = +1.72; 29.451, 7.90, −0.45, +0.39, +1.48, +1.70; 29.519, 7.89, −0.46, +0.37, +1.49, +1.71; 29.525, 7.89, −0.48, +0.36, +1.50, +1.72; 29.531, 7.89, −0.47, +0.36, +1.50, +1.72; 29.537, 7.89, −0.47, +0.37, +1.49, +1.71. Errors are likely to be ∼ 0.02 mag in all colors except U − B, where the error is likely to be 0.05 mag and too red.

P. Whitelock, South African Astronomical Observatory (SAAO), forwards the following infrared magnitudes obtained by A. Zijlstra (European Southern Observatory) on the 1.9-m SAAO telescope at Sutherland: Jan. 31.08 UT, J = 5.16, H = 5.46, K = 4.97, L = 4.08. J, H, and K are accurate to ± 0.03, and L to ± 0.05.

W. Liller, Viña del Mar, Chile, reports further CCD photometry: Jan. 30.14 UT, V = 7.47 (comparison stars HD 129614, HD 129738). J. Licandro, Montevideo, Uruguay, forwards the following visual magnitude estimate from Y. Fernandez: Jan. 31.8, 7.2.

Corrigendum. On IAUC 6130, line 17, for U − B = −0.44 ± 0.50, read U − B = −0.44 ± 0.05,

SUPERNOVA 1995A IN MCG +4-16-006 C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery

of an apparent supernova (V ∼ 18) at α = 6^{h}50^{m}04^{s}.63, δ = +26^{o}22′51′′.0 (equinox 2000.0); it was found on two technical pan films taken on Jan. 2.08 and 4.0 UT with the 0.9-m Schmidt telescope. The host galaxy is the southeastern member of the pair catalogued as MCG +4-16-006, and the supernova is located 0′′.5 east and 3′′.5 north of the host galaxy’s center; an OCA film on Jan. 28.0 UT shows that SN 1995A has faded by > 1 mag. It does not appear on the Palomar Sky Survey prints. A nearby star (V ∼ 17.5) is at position end figures 06^{s}.50, 20′′.7. The northwestern member of this pair of galaxies is located at end figures 03^{s}.72, 54′′.1.

1995 January 31 Daniel W. E. Green